Find the maximum height of the curve
y = 4 sinx - 3 cos x abive the x axis
1
Expert's answer
2021-09-28T11:13:31-0400

To find x at which y becomes maximum, find the first derivative by x and equate it to 0:


"y'=4\cos x+3\sin x =0,\\\space\\
\tan x=\frac 43,\\\space\\
x=53°.\\\space\\
y_\text{max}=4\sin53°-3\cos53°=1.39."